# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ 17 Guarujá precatório juros mora cálculos
Some interesting observations can be made from LABEL:tab:LIME_features. First, and as expected, we can see that many words are shared with the TF-IDF models (presented in LABEL:tab:importance_features). However, in the latter models, we noticed that the important features were almost exclusively words already present in the BP statement. In the case of BERT, LIME brings to light features which, although highly relevant to the identification of the legal theme covered by the BP, do not appear in its wording. This is the case of inquérito (inquiry/investigation) for BP 14, or reajuste (readjustment) for BP 37; we refer the reader to Sections 4.3 and 4.6 respectively for a legal analysis.
On the other hand, one sees that these important words are only detected by LIME in a limited number of documents. This is at most 57%, for the word algema (handcuffs) in BP 11. For BP 37, the most important word is isonomia (isonomy), which is detected in only 18.3% of documents. One might have expected these words to have been detected more frequently. This could explain the relatively low performance of the BERT models shown in LABEL:tab:similar_case_matching: the models are not sufficiently fit for the classification task.
4 Results and legal analysis
@@
This section is concerned with the application of our models, trained on the first dataset (documents citing a BP), to the analysis of the second dataset (whole collection of documents emitted by the High Court). From a legal point of view, we are interested in uncovering the juridical mechanisms behind the observed increase in cases citing a binding precedent, visualized in Fig. 1. Besides, from a Machine Learning perspective, we aim to verify the ability of the models to generalize their results, when applied to a larger dataset. We shall start with a general description of our methodology for empirical evaluation of binding precedents (Section 4.1), then give a detailed analysis of each BP (Sections 4.2, 4.3, 4.4, 4.5 and 4.6), and finally draw juridical conclusions (Section 4.7).
@@ 4.1 Description of our methodology
To evaluate the effect of a binding precedent on jurisprudence, we argue that one should not perform a mere reading of the cases citing that precedent, but also compare them to similar cases published before the precedent. This information can be visualized through the curve which associates, to each timestamp, the number of similar documents. In our context, we expect such a curve to behave as follows: the existence of a peak of documents just before the publication of the BP, followed by a steady decrease, until reaching a stable state, hopefully consisting of few documents. This behavior would indicate that the BP served its main purpose, which is to “settle” an increasingly contentious legal issue, clarifying for interested parties and the lower courts what the final position of STF is, thus reducing the likelihood of future conflicts involving the same issue.
If this behavior is not observed, as it is the case for the five precedents studied in this article, then one can add metadata information to the curve, to reveal the reasons for this deviant phenomenon. Namely, we will consider the type of process (e.g., habeas corpus or appeal), the state of provenance, the decision of the court (accepted, rejected or deferred), and the presence of certain words (specific to the legal subject considered).
@@
Generalization of the models
As discussed in Section 3.1, this article deals with Datasets #1 and #2, the former consisting of documents citing a binding precedent (29,743 documents), and the latter of all the documents in the categories “administrative law”, “criminal law” and “criminal procedure law” (615,262 documents). In Section 3.3, we trained the models on Dataset #1 and presented the results (see LABEL:tab:similar_case_matching). We will, in this section, apply these models on Dataset #2. In doing so, we face a significant problem: the models have been tuned to the first dataset, and are likely to give poor results on the second. As a matter of fact, we observed that the models tend to predict way more documents than intended, that is to say, they are underfitted.
To remedy this problem, we will not consider the binary output of the models, but rather their “probabilities”. As a native property of the models we considered, each document is associated with a likeliness of belonging to the positive class. In the case of SVM, this number is a distance; for the other models, it is a value between 0 and 1. Internally, a document is assigned the positive class if this value exceeds a certain threshold. To adapt the models to the new dataset, one should increase this threshold, discarding the false positives.
However, as already pointed out in Section 3.1, Dataset #2 does not come with indications of whether a document is related to a certain BP. Thus, this dataset cannot be used directly to tune the thresholds. Instead, we will take the following observation as a reference: if a document, issued after the publication of the BP, is related to the subject of this BP, then it probably cites the BP. In other words, we can take the annotations of Dataset #1 as ground truth, valid only after the publication of the BP. A well-tuned model is expected to follow faithfully the curve of citation of the BP.
In practice, we chose the new thresholds as the largest such a value that the recall of the predictions, after publication of the BP, is greater or equal to a fixed value $p\in[0,1]$. That is, we impose that at least a proportion $p$ of the groundtruth documents are detected. Note that the larger $p$ is, the more documents are detected. Specifically, we chose the values $p=0.9$ for BPs 11, 14 and 26, and $p=0.8$ for BPs 17 and 37. Although arbitrary, these values have been chosen in accordance with the models’ performance in LABEL:tab:similar_case_matching, and because they yield satisfactory results. We indicate in LABEL:tab:adjusted_probabilities, for each model and each BP, both the initial and the tuned value of the thresholds. For all cases, the tuned one is higher, indicating a model with fewer positive predictions. This will be also illustrated in the Figs. 3, 6, 10, 15 and 19, which represent the time series of estimated documents by the models, both for the initial thresholds (dashed curves) and the tuned thresholds (solid curves).
Table 5: For each model and each BP, the threshold over which a document is assigned the positive class is indicated. The first line refers to the initial threshold (computed from Dataset #1) and the second line to the tuned one (adapted to Dataset #2).
Model Threshold BP 11 BP 14 BP 17 BP 26 BP 37
TF-IDF+ Dataset #1 -0.7390 -0.6639 -0.9233 -0.6967 -0.7248
SVM Dataset #2 0.8867 0.9661 0.9101 0.9601 0.9428
TF-IDF+ Dataset #1 0.5 0.5 0.5 0.5 0.5
logistic Dataset #2 0.9967 0.9990 0.9978 0.9983 0.9649
TF-IDF+ Dataset #1 0.5 0.5 0.5 0.5 0.5
forest Dataset #2 0.59 0.662 0.611 0.642 0.538
LSTM Dataset #1 0.9999991 0.9970 0.99992 0.9997 0.9848
Dataset #2 0.9999999 0.9999 0.99995 0.9999 0.9999
BERT Dataset #1 0.8431 0.9676 0.7256 0.1393 0.9543
Dataset #2 0.9906 0.9928 0.9995 0.9955 0.9998

### v3

@@ 3.4 Explainability
The table offers several interesting observations. First, and as expected, one sees that many words are shared with the TF-IDF models (presented in Table 3). However, in the latter models, we noticed that the important features were almost exclusively words already present in the BP statement. As a matter of fact, among the five words that were not, two were also detected by LIME (criminal and inquérito). In the case of Longformer, LIME brings to light features that, although highly relevant to the identification of the legal theme covered by the BP, do not appear in its wording.
This is the case for BP 11: the words despacho (procedural ruling) and reclamação (appeal/complaint) indicate that the model Longformer tends to favor documents whose type is Reclamação, and which have not been judged on their merits, but deferred for procedural reasons. This is precisely what will be observed in the dedicated legal section, and explains the recent use of this BP (see Section 5.2).
BP 26, presenting the words SP and geral, is also noteworthy. The first refers to the State of São Paulo, which, as studied in Section 5.5, is precisely where most of the debate surrounding the BP is taking place. Equally relevant, the word geral is found in Procurador-Geral or Procuradoria-Geral (General Prosecutor, General Prosecutor’s Office), present in 58.9% of documents citing BP 26. In fact, prosecutors play a particular role in the context of BP 26, since it is they who may request a criminological examination, an issue that has become central to the abovementioned debate.
Finally, LIME detects two words in the last BP that are not part of its statement: 2015 and magistratura (Judiciary). As we will see in Section 5.6, BP 37 is paved with requests by different groups of litigants; in particular, major peaks include one in 2015 (from temporary professors) and another in 2017 (from Judiciary members).
4 Generalization on Dataset #2
In the previous section, we trained models on the first dataset (documents citing a BP). Now we’d like to assess the quality of their predictions when applied to the second dataset (the whole collection of documents emitted by the Supreme Court in the selected branches). That is to say, we seek to verify their ability to generalize their results when applied to a larger dataset. First of all, a fine-tuning step is required, as detailed in Section 4.1. Then, we will check the quality of the predicted documents by three means: the study of the time series of predicted documents, the correlation between the words of these documents, and a manual reading of a few samples. This analysis will be carried out precedent by precedent, in Section 4.2.
4.1 Evaluating the generalization of the models
Fine-tuning
When applying the models, trained on Dataset #1 (29,743 documents) to Dataset #2 (634,068 documents), we face a significant problem: the models have been tuned to the first dataset, and can potentially be not adapted to the second. In fact, we observed that they tend to predict way more documents than intended, i.e., they are underfitted.
To remedy this problem, we will not consider the binary output of the models, but rather their “probabilities”. As described in Section 3.3, as a native property of the models we considered, each document is associated with a likeliness of belonging to the positive class, through a value between 0 and 1. Internally, a document is assigned the positive class if this value exceeds the threshold of $0.5$. To adapt the models to the new dataset, one should increase this threshold, to discard the false positives.
However, as already pointed out in Section 3.1, Dataset #2 does not come with indications of whether a document falls within the scope of a certain BP. Thus, this dataset cannot be used directly to tune the thresholds. Instead, we will take the following observation as a reference: if a document, issued after the publication of the BP, is detected by the models, then it probably cites the BP. In other words, we can take the labels of Dataset #1 as groundtruth, valid only after the publication of the BP. A well-tuned model is expected to follow faithfully the curve of citations post-BP.
In practice, we chose the new thresholds as the largest value such that the recall of the predictions, after the publication of the BP, is greater or equal to a fixed value $p\in[0,1]$. That is, we impose that at least a proportion $p$ of the groundtruth documents are detected. Note that the larger $p$ is, the more documents are detected. Specifically, we chose the values $p=0.9$ for BPs 11, 17, and 37, and $p=0.7$ for BPs 14 and 26. Although arbitrary, these values have been chosen in accordance with the models’ performance in Table 2, and because they yield satisfactory results. We indicate in Table 5, for each model and each BP, the tuned value of the thresholds. For all cases, the tuned threshold is higher than $0.5$, the native threshold, implying a model with fewer positive predictions.
Model BP 11 BP 14 BP 17 BP 26 BP 37
TF-IDF+SVM 0.999 0.995 0.999 0.999 0.987
TF-IDF+logistic 0.998 0.999 0.998 0.999 0.951
TF-IDF+forest 0.7 0.869 0.697 0.893 0.74
LSTM 0.998 0.998 0.997 0.999 0.995
Longformer 0.615 0.999 0.999 1 0.997
Table 5: For each model and each BP, the threshold over which a document is assigned the positive class. The thresholds refer to those tuned on Dataset #2 and are given with three decimal places.
Time series of similar cases
Our main tool for visualizing model predictions on Dataset #2 will be time series, displayed as curves. More precisely, we will represent the quantity of documents predicted by each model as a function of their publication date, in the form of a histogram, which is then interpolated by a curve. This representation has the advantage of revealing considerable information. Firstly, from a purely methodological point of view, we can assess the predictive quality of the models: as explained above, we expect a fine-tuned model to follow closely the curve of documents citing the BP considered (we call it the groundtruth curve), at least after the date of publication of this BP. Before the publication date, other phenomena can be observed. We expect certain legal events to induce peaks in these curves, which can then be identified visually, and investigated further by reading the respective documents, or by adding additional metadata. These curves also show the overall trends in BP citations, which we will use for our legal analysis in Section 5. Lastly, by plotting the prediction curves of several models on the same graph, we obtain greater confidence in the legal events detected. Indeed, if several models present a peak at the same time, it may well be a quantity of documents effectively linked to the content of the BP. On the other hand, if a trend is exhibited by only one model, we might suspect that it is the result of poor training.
We specify that not all the documents detected by a model will be considered, but only those belonging to the branch of the BP under consideration — “Administrative Law” for BPs 17 and 37, and “Criminal Law” or “Criminal Procedure Law” for BPs 11, 14, and 26. Indeed, since the scope of a BP is reduced to a precise branch, as described in Section 3.1, documents predicted outside this category would be false positives, and therefore irrelevant. We have observed that, in practice, this adjustment brings only a minor change in the prediction curves.
Analyzing these curves, and assessing whether the predicted documents are relevant, will be the content of Section 4.2. In particular, the time series will be given in Figs. 7, 9, 12, 14 and 15 for BP 11, 14, 17, 26 and 37, respectively.
Correlations of words
4.2 Manual evaluation
In this section, we will analyze the predictions of our models precedent by precedent, in order to check that they are correct, i.e., that the predicted documents are indeed related to the content of the precedent. As we will see, some sets of predictions will need to be discarded. This section is therefore intended as a validation phase for the models, enabling the final legal analysis of the precedents in Section 5. To reserve legal considerations for the final section, we will aim to minimize here references to the legal content of the documents and instead focus primarily on analyzing them through the lens of time series, word correlations, and regular expressions.
Binding Precedent 11
We show in Fig. 7 the predictions of the five models on our whole collection (Dataset #2), as well as a regex search for documents containing the words algemas or algemado (handcuffs or handcuffed). As explained in Section 4.1, the thresholds of the models, trained on Dataset #1, are not adapted for Dataset #2: way too many documents are predicted. Therefore, we choose new thresholds with the following rule: the value such that 90% of the documents citing BP 11 are predicted. In the figure, the initial and the fine-tuned prediction curves are visualized as dashed or solid, respectively.
Models’ predictions for BP 11 Refer to caption
Figure 7: Number of documents predicted by each model for BP 11 in Dataset #2, represented as a histogram (window length of one year) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and Dataset #2 (solid), as given in Table 5. Two views are given, the second one zooming in on the ordinate axis.
As it turns out, there are almost no documents, before the publication of BP 11, containing the words of the regex query. This phenomenon is also illustrated by the models TF-IDF+SVM, TF-IDF+logistic, TF-IDF+random forest, and Longformer, which closely follow the groundtruth curve. As a matter of fact, among the 1056 documents of Dataset #2 containing the words algemas or algemado, respectively 656, 679, 656, and 814 were predicted by the models, accounting for $88.3\%$, $82.9\%$, $93.6\%$, and $91.2\%$ of their predictions. In these documents, these words are often used in conjunction with uso and nulidade (use and nullity), also present in the official wording of BP 11. These observations illustrate the fact that the documents processed by STF have only begun to mention “handcuffs” since the publication of BP 11. This is an interesting legal fact, which we will explore in Section 5.2 dedicated to legal analysis.
The model LSTM, however, estimates thousands of documents. Only 4.4% of its predictions contain the words algemas or algemado. A manual inspection shows that this model tends to detect documents merely containing the words uso or nulidade, independently of the surrounding context. For instance, among the 52,465 documents containing the word uso, 9586 are associated with the positive class, accounting for 60.4% of the predicted documents. This suggests that LSTM is biased toward detecting this word, regardless of its role in the conjunction uso de algemas (use of handcuffs).
To analyze this phenomenon further, let us consider the TF-IDF’s most relevant words, obtained in Table 3: algemado, algemas, audiência, nulidade, and uso (handcuffed, handcuffs, court hearing, nullity, use). We give in Fig. 8 the correlations between these five words when restricting the dataset to the documents explicitly citing BP 11, as well as the documents estimated by the models. As explained in LABEL:{subsec:evaluating_generalization}, a well-fitted model is expected to show similar correlations as the groundtruth. The first image shows a high correlation between uso and algemas, indicating that they mainly belong to the set uso de algemas. Besides, only LSTM does not reflect this behavior. This observation accounts for the fact that this model, by not taking into account the correlation between these words, shows poor results when used on the larger dataset.
Figure 8: Correlations between the selected relevant words, when restricted to documents citing BP 11 (“Groundtruth”) or documents predicted by one of our models. Note that, as discussed in LABEL:{subsec:evaluating_generalization}, correlations are calculated by seeing the words as binary random variables, hence correlation values lie between 0 and 1.
Finally, we turn our attention to the jump in regex predictions, after 2016, which is not part of the groundtruth (i.e., the documents do not cite BP 11), seen in Fig. 7. A manual inspection of these documents reveals that, although they contain the words algemas or algemado, are not directly related to the topic of BP 11. As a matter of fact, these documents mention the (non-binding) Precedent 279  [^9] Precedent 279: “For the mere reexamination of evidence, an extraordinary appeal (ARE) is not admissible” https://portal.stf.jus.br/jurisprudencia/sumariosumulas.asp?base=30&sumula=2174 (Súmula 279). In addition to Precedent 279, the documents cite prior decisions based on it, and in particular ARE 965.920101010ARE 965.920 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=11766746, which contains the word algemas. Consequently, the regex search returned these documents. However, this detection is merely coincidental: the question of handcuffs is not at the core of the documents. This illustrates the limitations of a simple regex search, as well as the importance of verifying, by reading, the predictions obtained. In particular, we will not use these predictions to conduct our legal analysis. To illustrate the issue more specifically, we highlight that ARE 965.920 is a case where BP 11 and Precedent 279 intersect. The decision’s rationale indicates that demonstrating the improper use of handcuffs (the subject of BP 11) would require reanalyzing evidence, which falls under the scope of Precedent 279. Consequently, the non-binding precedent was invoked to justify denying the appeal.
Binding Precedent 14
The prediction curves for BP 14 are displayed in Fig. 9. We remind the reader that, for this precedent, we tuned the models’ thresholds as those such that 70% of the documents citing BP 14 are predicted, and not 90%, as is the case for BPs 11, 17, and 37 (see Table 5). A quick analysis of the predicted documents indicates that, among the five models considered, only TF-IDF+random forest achieved the expected result, by faithfully following the groundtruth curve. This is also the case for the regex search of acesso aos elementos/autos/documentos (access to elements, records, or documents).
Models’ predictions for BP 14 Refer to caption
Figure 9: Number of documents predicted by each model for BP 14 in Dataset #2, represented as a histogram (window length of one year) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and Dataset #2 (solid), as given in Table 5. Two views are given, the second one zooming in on the ordinate axis.
Interestingly, all the other models peaked in 2017/2018. To investigate why they estimate many documents, we will not consider the correlations, as we did for BP 11 (in Fig. 8), but simply the frequencies of the most relevant words, represented in Fig. 10. Almost all groundtruth documents cite acesso aos autos (access to records). Similarly, 95.8% of TF-IDF+random forest’s predictions mention autos, of which 92.1% also mention acesso. Besides, among TF-IDF+SVM, TF-IDF+logistic, LSTM and Longformer’s predictions, respectively, 95%, 95.5%, 90.8%, and 86.4% mention autos, while only 64.6%, 66.3%, 48.7%, and 41.7% of them conjointly contain acesso. This suggests some underfitting of the models, which are unable to identify the two words in combination.
At first sight, a manual inspection of a sample of the detected documents, excluding elements of the groundtruth class, raises questions about the qualities of these predictions. Although certain decisions [^11] Such as Inq 4.260 https://portal.stf.jus.br/processos/downloadPeca.asp?id=313917544&ext=.pdf and RHC 133.298 https://portal.stf.jus.br/processos/downloadPeca.asp?id=315014507&ext=.pdf are on topics regarding investigations (specifically, on the initiation of an official investigation and the transcription of some telephonic tapping record), neither of them concerns topics related to access of investigation material by the investigated part, which is the subject of the BP. The addition of metadata, as in Fig. 11, shows that this peak mainly consists of cases from Distrito Federal that also mention sigilo (secrecy/confidentiality). This suggests that these cases are complaints filed by attorneys representing individuals investigated or condemned for white-collar crimes, most relevantly, by attorneys representing politicians or individuals of political interest. It is, therefore, unsurprising that this increase in predictions after 2012 partly matches the interval when Operação Lava Jato took place, a joint cooperative investigation aiming at politicians accused of corruption-related crimes. We continue this discussion, from a legal perspective, in Section 5.3.
Additional metadata for documents predicted by Longformer for BP 14 Refer to caption
Figure 11: Top: State to which the cases predicted by Longformer for BP 14 belong, with label “UNKNOWN” when the metadata is unavailable. Bottom: Regex search for the presence of the words sigilo (secrecy/confidentiality) and inquérito (inquiry/investigation). In both graphs, the bins have a length of 6 months.
Binding Precedent 17
Fig. 12 presents the predictions of the five models for BP 17, together with a regex search for precatório and juros de mora (court orders, late payment interest). Some common trends stand out: a peak between 1995-1998 (by regex, TF-IDF+SVM, and logistic regression), one between 2002-2005 (by LSTM and Longformer, but also TF-IDF+SVM and logistic), one in 2007-2009 (by regex and Longformer), and a last peak in 2014-2016 (after the publication of the BP, by regex and LSTM), as well as an increasing tendency starting from 2011, predicted by all the models. Besides, LSTM and Longformer individually detect two other peaks: in 1998 and 1996, respectively.
Models’ predictions for BP 17 Refer to caption
Figure 12: Number of documents predicted by each model for BP 17 in Dataset #2, represented as a histogram (window length of one year) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and Dataset #2 (solid), as given in Table 5. Two views are given, the second one zooming in on the ordinate axis.
As with all the other binding precedents, BP 17 had random forest for the best model, in the sense that it predicts better the groundtruth curve, after the publication of the precedent. However, in the context of BP 17, this model is not of interest: the many legal events that took place before the publication of the BP, and which will help us to analyze it in Section 5.4, are not captured by TF-IDF+random forest.
The difference between the model’s predictions is directly explained by the words correlation in Fig. 13: except TF-IDF+random forest (and also LSTM to a lesser extent), none of the models accurately captured the correlation of precatório (court order payment) with the others words (mora, late payment). In particular, the most important peak, between 2002-2005, is principally composed of documents mentioning “court order” but not “late payment”. This explains the absence of regex’s predictions in this period, once more highlighting the limitations of a simple regex search. This peak, even though not referring to cases directly linked to the theme of BP 17, contains, as we will analyze, relevant information.
Another interesting observation is the significant correlation, in most of the models, between the terms moratório and juros/mora. The term moratório (moratorium) can, indeed, be used as a synonym for mora (late). However, it seems that the models have imposed that predicted documents, when they contain the latter, must contain the former. This would explain why regex’s peak in 2007, mainly composed of documents citing juros de mora but not moratório, has only been detected by Longformer.
As it turns out, Longformer’s peak in 1996 consists solely of correct predictions, as confirmed by manual reading. Their analysis (involving Article 33 of the Act of Transitory Constitutional Dispositions) will be carried out in Section 5.4 (Fig. 19). Interestingly, although 120 of these documents contained both the terms precatório and juros de mora, being therefore also detected by regex, many others had only the word precatórios and reference to Article 33. In this sense, the neural network model could identify cases strongly related to the BP, but for which our regex terms would be blind.
LSTM’s 1998 peak is more challenging to understand. All documents correspond to appeals promptly rejected by the STF based on procedural arguments, with little or no analysis of the materiality of the cases. In addition, of the 221 documents detected by LSTM between 1997 and 1999, only 10 contained the term precatório and two contained mora, although, surprisingly, 31 contained terms related to monetary corrections of inflation, which, in principle, has little to do with the BP. Because all these documents were particularly short (with an average of 91 tokens), it is possible to hypothesize that LSTM was, for some reason, biased towards short decisions for this particular BP, although we do not further explore this possibility here.
To conduct our legal analysis, we will examine in greater depth the predictions offered by Longformer and regex; the addition of metadata will help reveal the underlying legal context (see Figs. 19 and 20, respectively). These two models cover all the peaks observed in Fig. 12, with the exception of LSTM in 1998, deemed irrelevant. More specifically, we will study the peaks detected separately by Longformer and regex in 2002-2005 and 2014-2016, respectively, and those detected jointly in 1996 (the latter’s predictions are a subset of the former) and 2007-2009.
Binding Precedent 26
Fig. 14 shows the predictions of the five models, as well as a regex search for the words exame criminológico and progressão de regime (criminological examination, regime progression). Except for random forest, they all present a peak in 2006-2008. For all models, one also notes an increasing tendency starting from 2014. Both observations correspond to legal events specific to BP 26 and will be discussed in Section 5.5.
Models’ predictions for BP 26 Refer to caption
Figure 14: Number of documents predicted by each model for BP 26 in Dataset #2, represented as a histogram (window length of one year) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and Dataset #2 (solid), as given in Table 5. Two views are given, the second one zooming in on the ordinate axis.
As far as the 2006-2008 peak is concerned, the regex search finds 659 documents in this period. A manual inspection of these documents shows that they are indeed related to the theme of BP 26, hence are considered correct predictions of the models.
To explain why TF-IDF+random forest did not detect this peak, let us consider two of its most important features (not presented in Table 3): the words realização (realization) and subjetivos (subjective) are present in respectively 86.6% and 78.1% of its 529 predicted documents (between 2012-2019). In opposition, they are only present in 60.4% and 60.8% of the 659 documents detected by regex between 2006-2008. This suggests that this TF-IDF model has learned unnecessary restrictions, imposing the presence of words realização and subjetivos, resulting in too few predicted documents. In our legal analysis, we will not consider the predictions of TF-IDF+random forest, the other models being deemed to contain more valuable information.
Binding Precedent 37
We show in Fig. 15 the predictions of the five models, together with a regex search for the words isonomia, vencimentos and servidores públicos. We also add a search for Súmula 339 (Precedent 339), which, as we will see in Section 5.6, is responsible for the creation of BP 37, and can therefore be used as a groundtruth before the precedent’s publication date. As demonstrated in the bottom part of the figure, TF-IDF+random forest presented the best results, along with regex, since they closely follow the groundtruth curve after the BP’s publication.
Models’ predictions for BP 37 Refer to caption
Figure 15: Number of documents predicted by each model for BP 37 in Dataset #2, represented as a histogram and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and Dataset #2 (solid), as given in Table 5. Two views are given, the second one zooming in on the ordinate axis. In addition, we represent a regex query for Súmula 339. Unlike the other BPs, the bins in this figure are six months long, to observe the peaks more precisely.
One observes a few common trends in the figure. Before the publication of the BP, the model Longformer, as well as regex, show a clear peak in 2005. These documents indeed contain words that characterize BP 37. More specifically, a manual inspection shows that the documents involve a request for salary readjustment, requested by military personnel, based on Law 8.622/1993121212Law 8.622/1993 https://www.planalto.gov.br/ccivil_03/LEIS/L8622.htm, which the STF addressed on the grounds of Article 37, Section X of the Constitution. This observation is confirmed by Fig. 16, which shows the presence of the words 8.622/93 and Artigo 37 (Article 37) in Longformer’s predictions, found by regex. This event is part of the history of Precedent 339, studied in the dedicated juridical section (see the discussion of Fig. 23).
Additional metadata for documents predicted by Longformer for BP 37 Refer to caption
Figure 16: Top: Type of legal process used in the documents predicted by the model Longformer for BP 37, among RE (Recurso Extraordinário), Rcl (Reclamação), and ARE (Recurso Extraordinário com Agravo). Bottom: Documents containing the words artigo 37 (Article 37) or 8.622/93 (the law). Bins are six months long.
Besides, after the publication of BP 37, three peaks in the number of predictions can be observed (in late 2015, 2016, and 2017), reflected by all the models. As it turns out, they can be mapped to specific juridical events, as studied further (see Fig. 22).
Last, to understand the low number of predictions of the models before 2012 compared to the documents citing Precedent 339, one can refer to the important words of TF-IDF, collected in Table 3. In particular, isonomia has the greatest importance, hence we expect that a large portion of the predicted documents contain this word. It is indeed the case, at a rate of 77.2%, 76.3%, and 89.6% for SVM, logistic regression, and random forest. Comparatively, only 65.2% of the documents citing Precedent 339 mention this word. This suggests that the models are biased towards estimating principally documents employing isonomia, which is not representative of all the documents similar to BP 37. A similar observation can be made for the word aumentar, explaining the small number of documents predicted.
5 Legal analysis
@@
For the last step of our program, we will use the models, trained on Dataset #1 in Section 3, and whose predictions on Dataset #2 have been validated in Section 4, to help uncover the juridical mechanisms behind the observed increase in cases citing a binding precedent (visualized in Fig. 1). We shall start with a general description of our methodology for empirical evaluation of BP (Section 5.1), then give a detailed analysis of each BP (Sections 5.2, 5.3, 5.4, 5.5 and 5.6), and finally draw juridical conclusions (Section 5.7).
@@ 5.1 Description of our methodology
To evaluate the effect of a binding precedent on jurisprudence, we argue that one should not perform a mere reading of the cases citing that precedent, but also compare them to similar cases published before the precedent. As already presented in Section 4.1, this information can be visualized through the curve which associates, to each timestamp, the number of similar documents. In our context, we expect such a curve to behave as follows: the existence of a peak of documents just before the publication of the BP, followed by a steady decrease, until reaching a stable state, hopefully consisting of few documents. This behavior would indicate that the BP served its main purpose, which is to “settle” an increasingly contentious legal issue, clarifying for interested parties and the lower courts what the final position of the STF is, thus reducing the likelihood of future conflicts involving the same issue.
If this behavior is not observed, as is the case for the five precedents studied in this article, then one can add metadata information to the curve, to reveal the reasons for this deviant phenomenon. Namely, we will consider the type of process (e.g., habeas corpus or appeal), the state of provenance, the decision of the court (accepted or rejected), the category of juridical reasoning (based on the merits of the case or due to procedural flaws), and the presence of certain words (specific to the legal subject considered).